40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 12 18:08:54 2020, gautam, Update, Electronics, Number of the beast IMG_8710.JPG
    Reply  Wed Aug 12 20:15:42 2020, Koji, Update, Electronics, Number of the beast 
    Reply  Tue Sep 15 12:20:36 2020, gautam, Update, Electronics, SR785 repaired dividerTF.pdf
Message ID: 15571     Entry time: Tue Sep 15 12:20:36 2020     In reply to: 15517
Author: gautam 
Type: Update 
Category: Electronics 
Subject: SR785 repaired 

The unit was repaired and returned to the 40m. Now, with a DMM, I measure a DC offset value that is ~1% of the AC signal amplitude. I measured the TF of a simple 1/20 voltage divider and it looks fine. In FFT mode, the high frequency noise floor levels out around 5-7nV/rtHz when the input is terminated in 50 ohms.

I will upload the repair documents to the wiki.

Quote:

The "source" output of the SR785 has a DC offset of -6.66 V. I couldn't make this up.

Attachment 1: dividerTF.pdf  24 kB  Uploaded Tue Sep 15 13:33:06 2020  | Hide | Hide all
dividerTF.pdf
ELOG V3.1.3-